How to calculate normal distribution between 60 and 70 with a mean 65 and SD of 5?


1
Expert's answer
2021-03-07T16:57:57-0500

Let "X\sim N(\mu, \sigma^2)." Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)"

Given "\mu=65, \sigma=5"


"P(60\leq X\leq70)=P(X\leq 70)-P(X<60)"

"=P(Z\leq \dfrac{70-65}{5})-P(Z<\dfrac{60-65}{5})"

"=P(Z\leq 1)-P(Z<-1)"

"\approx 0.841345-0.15866\approx0.6827"

The probability is 0.6827.